# Stochastic Geometry and Dynamical System Analysis of Walker Satellite Constellations

### v2

@@ III-A Invariant Probability Measure
Let $\overline{\cal Q}$ denote the product of the two uniform distributions on the product space: $S=[0,\frac{2\pi}{N_{o}}]\times[0,\frac{2\pi}{N_{o}}]$. Let $\overline{\cal Q}_{\Psi}$ denote the distribution of $\Psi$ at time zero when $(\bar{\theta},\bar{\omega})$ has the distribution $\overline{\cal Q}$.
@@
The probability measure $\overline{\cal Q}$, on $S$, is invariant for $R$ and the probability measure $\overline{\cal Q}_{\Psi}$, on $\mathcal{N}_{r}$, is invariant for $\Psi$.
@@
It is enough to prove the first statement. Let $\mathbb{T}[0,a]\times[0,b]$ denote the torus on $[0,a]\times[0,b]$, with $a,b$ positive real numbers. Let $\theta=\bar{\theta}/{2\pi}$ and $\omega=\bar{\omega}/{2\pi}$. Let also $\theta(t)=\bar{\theta}(t)/{2\pi}$ and $\omega(t)=\bar{\omega}(t)/{2\pi}$. Denote by $g$ the linear map that transforms $(\theta(t),\omega(t))$ into $(\bar{\theta}(t),\bar{\omega}(t))$. The associated dynamics w.r.t. time is $\alpha(t)({\theta},{\omega})=({\theta}(0)-v_{{\theta}}t\mod 1,{\omega}(0)+v_{{\omega}}t\mod 1)$, where $v_{{\theta}}=\bar{v}_{{\theta}}/{2\pi}$ and $v_{{\omega}}=\bar{v}_{{\omega}}/{2\pi}$. Then, $\overline{\cal Q}$ on $\bar{\mathbb{T}}([0,{2\pi}],[0,{2\pi}])$ is invariant for $\bar{\alpha}(t)(\bar{\theta},\bar{\omega})=(\bar{\theta}-\bar{v}_{\theta}t{\mod 2\pi},\bar{\omega}+\bar{v}_{\omega}t{\mod 2\pi})$ if and only if $\cal Q$ on $\mathbb{T}([0,1],[0,1])$, namely the product of the two uniform distributions on $[0,1]$, is left invariant for $\alpha(t)$. But this fact is well known for all nonzero values of $(v_{{\theta}},v_{{\omega}})$ [13]. ∎
@@ III-B Ergodicity and Periodicity
If the ratio $\rho$ is irrational, then the dynamical system $R$ (resp. $\Psi$) on state space $S$ (resp. $\mathcal{N}_{r}$) is ergodic w.r.t. $\overline{\cal Q}$ (resp. $\overline{\cal Q}_{\Psi}$). If the ratio $\rho$ is rational, then $R$ (resp. $\Psi$) is periodic.
@@
It is enough to prove the statement concerning $R$. We use the notation in the proof of Theorem 1. Let $\hat{\rho}$ denote the normalized ratio ${v_{\theta}}/{v_{\omega}}$. If $\hat{\rho}$ is irrational, then the dynamical system $\alpha(t)$ on $\mathbb{T}([0,1],[0,1])$ is minimal and ergodic w.r.t. $\cal Q$, namely the product of two uniform distributions on $[0,1]$ (see e.g. [14, Section 6.5.2] and [15, Theorem 2.1]). If $\hat{\rho}$ is rational (which is equivalent to having the ratio $\rho$ rational), then $\alpha(t)$ is periodic on $\mathbb{T}([0,1],[0,1])$. Hence, it is not ergodic w.r.t. $\cal Q$ (see [14, Section 6.5.2] and [15, Theorem 2.1]). ∎
@@
When the ratio $\rho$ is irrational, we have hence
$\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}h(\Psi(x,t))\mathrm{d}t=\int h(\psi){\overline{\cal Q}}_{\Psi}(\mathrm{d}\psi),$ (8)
for all functions $h:{\mathcal{N}}_{r}\to\mathbb{R}$ such that $\int|h(\psi|){\overline{\cal Q}}_{\Psi}(\mathrm{d}\psi)<\infty$, where the limit holds for $\overline{\cal Q}$ almost all $x$. In other words, any time average is given by the expectation w.r.t. the time-invariant distribution ${\overline{\cal Q}}_{\Psi}$ of $\Psi$.
In contrast, if $\rho$ is rational, the system is then periodic. The empirically-obtained time average of what is observed by a given Earth user does not match averages over the stationary distribution $\overline{\mathcal{Q}}_{\Psi}$. Since the base point process $\Psi(x,t)$ for the satellite patterns is periodic in the Earth referential, any process determined by it is also periodic. For instance, suppose a user is at the fixed position $(x_{1},y_{1},z_{1})$ on Earth. Then, the satellite pattern seen from this user is periodic too. These periodic patterns observed however depend on the user location $(x_{1},y_{1},z_{1})$. For another user at $(x_{2},y_{2},z_{2})$, the satellite patterns seen are also periodic but with a structure possibly different from the one at $(x_{1},y_{1},z_{1})$.
@@
Since the altitude basically determines the satellite speed, the designer can choose an altitude making the constellation periodic to fixed Earth observers. However, an unaware designer who would pick a real number, say between $350$ km to $1500$ km, would get an irrational ratio for almost all choices w.r.t. the Lebesgue measure on this interval since the rational numbers have Lebesgue measure zero in the given interval. Therefore, based on Theorems 1 and 2, one can claim that designers unaware of this dynamical system property have ”almost surely” built constellations which are aperiodic and ergodic, and where the empirical averages over long time converge to the expectation over the invariant distribution $\overline{\mathcal{Q}}_{\Psi}$ studied in the present paper.
Below, we discuss the significance of the invariant measure (in all cases including the periodic case).
@@
Assume that the designer selects the angular parameters $\bar{\theta}$ and $\bar{\omega}$ as random variables sampled independently, with each of them being uniform on the relevant interval. Then, when deconditioning the position of the satellites of the constellation w.r.t. these two random variables, one gets a distribution for the satellite positions which is invariant over time, and which is precisely the distribution ${\overline{\cal Q}}_{\Psi}$ used in the present paper and leads to the specific distributions that we evaluate below for the metrics of interest. By the strong law of large numbers, the distributions computed here can hence also be seen as the empirical averages of the distributions evaluated over a large collection of samples of the constellation obtained from i.i.d. samples of these two angles. This observation is valid in all cases including the periodic case. Of course, in the (most likely) case of ergodicity, the metrics evaluated below are also time averages in the first place.
@@ IV-A Distance to Nearest Satellite
Let $\operatorname{\mathbb{S}}$ be the sphere of radius $r.$ For $r-e<d<\sqrt{r^{2}-e^{2}},$ we define the spherical cap w.r.t. the typical user $u$ and the distance $d$ as follows: $\operatorname{\mathbb{S}}_{u}(d)=\{(x,y,z)\in\operatorname{\mathbb{S}}|\|(x,y,z)-\vec{u}\|\leq d\}$. In particular, $\operatorname{\mathbb{S}}_{u}(\sqrt{r^{2}-e^{2}})$ is the visible spherical cap from the typical user and we simply denoted it by $\operatorname{\mathbb{S}}_{u}.$ Let $D(l_{u})$ be the distance from the typical user at latitude $l_{u}$ to its nearest satellite. $D(l_{u})=\infty$ if no satellite is visible from the typical user.
@@
$k\int_{0}^{\frac{2\pi}{N_{o}}}\int_{0}^{\frac{2\pi}{N_{s}}}\!\!\mathbbm{1}\left\{\frac{r^{2}+e^{2}-d^{2}}{2re}>\max\frac{\vec{X}_{i,j}\cdot\vec{u}}{\|\vec{X}_{i,j}\|\|\vec{u}\|}\right\}\mathop{}\!\mathrm{d}\omega\mathop{}\!\mathrm{d}\theta,$
where $\vec{X}_{i,j}$ is in Eqs. (4), (5), and (6). $k=N_{o}N_{s}/(4\pi^{2}).$
@@
Let $\mathbbm{1}\{x\}$ be the indicator function which takes value one if $x$ occurs and zero otherwise. Conditionally on $\bar{\theta},\bar{\omega},$ the distance from the typical user to its nearest satellite is
@@
where $(i,j)\in([N_{o}],[N_{s}])$ stands for $i=1,...,N_{o}$ and $j=1,...,N_{s}.$ To get this, we use the fact that all points of such orbits are not in the cap if and only if the central angle between the user and the satellite closest to the user is greater than $\kappa,$ where $\kappa=\arccos\left(\frac{e^{2}+r^{2}-d^{2}}{2re}\right)$ is the central angle from the user to the rim of the cap $\operatorname{\mathbb{S}}_{u}(d)$. Let $k={N_{o}N_{s}}/{(4\pi^{2})}$. Then,
@@
$=k\int_{0}^{\frac{2\pi}{N_{o}}}\int_{0}^{\frac{2\pi}{N_{s}}}\mathbbm{1}\left\{\cos(\kappa)>\max_{(i,j)}\frac{\vec{X}_{i,j}\cdot\vec{u}}{\|X_{i,j}\|\|u\|}\right\}\mathop{}\!\mathrm{d}\bar{\omega}\mathop{}\!\mathrm{d}\bar{\theta}$
$=k\int_{0}^{\frac{2\pi}{N_{o}}}\!\int_{0}^{\frac{2\pi}{N_{s}}}\!\!\mathbbm{1}\left\{\frac{r^{2}+e^{2}-d^{2}}{2re}>\max_{(i,j)}\frac{\vec{X}_{i,j}\cdot\vec{u}}{\|\vec{X}_{i,j}\|\|\vec{u}\|}\right\}\!\mathop{}\!\mathrm{d}\bar{\omega}\mathop{}\!\mathrm{d}\bar{\theta},$
@@
Thanks to the deterministic structure of the deployment of satellites, for latitudes $l_{u}$ smaller than a threshold $L$, for $N_{o}$ and $N_{s}$ large enough, there exists a constant $d_{c}(l)\leq\sqrt{r^{2}-e^{2}}$ such that $\operatorname{\mathbf{P}}[D(l_{u})>d]=1$ for all $d>d_{c}(l_{u})$. For instance, when $N_{o}=30$ and $N_{s}=50,$ we have $\operatorname{\mathbf{P}}(D({l_{u}})>d_{c}(l_{u}))=0$ for $d_{c}(l)\approxeq 775$km, e.g., the association distance is for sure less than $775$ km. This information is useful to establish the performance lower bound since it directly links to the maximum path loss of the downlink communications. The existence of this critical distance comes from the fact that proposed network features regularly placed satellites with a rigid structure, in contrast to other models based on binomial, Poisson, or Cox point processes [2, 3, 4, 5, 6, 7, 8, 9, 10]. Finding an explicit expression for $d_{c}(l)$ is left for future work.
@@
where the left-hand side is the time average of the association distance for a user and the right-hand side is its expectation w.r.t. the invariant distribution, that we obtained in Theorem 3.
@@
Fig. 2 shows the mean of the total interference. The simulation results confirm the accuracy of the derived formula. Note, the average total interference is the maximal for users at latitude $30\degree$. This indicates that the users around latitude $30\degree$ experience more total interference than the users at other latitudes, on average.
@@
[…]
@@
with $\tilde{c}_{i,j}\!=\!\mathbbm{1}\!\left\{\frac{\vec{X}_{i,j}\cdot\vec{u}}{\|\vec{X}_{i,j}\|\|\vec{u}\|}\geqq\frac{e}{r},\frac{\vec{X}_{i,j}\cdot\vec{u}}{\|\vec{X}_{i,j}\|\|\vec{u}\|}\neq\max\left(\frac{\vec{X}_{i,j}\cdot\vec{u}}{\|\vec{X}_{i,j}\|\|\vec{u}\|}\right)\!\right\}$ and
@@
As in Theorem 4, Theorem 5 yields a simpler expression for a Rayleigh fading scenario where $\mkern 1.5mu\overline{\mkern-1.5muF\mkern-1.5mu}\mkern 1.5mu_{H}(x)=e^{-x}$. Fig. 3 illustrates the coverage probability for the Rayleigh fading. From the figure, we confirm the accuracy of the derived formula.
The practical implication of employing the proposed geometric model lies in the fact that the derived formula computes the coverage probability of the typical user at a given latitude and it is equal to the coverage probability of all users at the same latitude. In addition, in the ergodic case in Theorem 2, the expression theoretically matches the empirically-obtained time average of the coverage probability of any user at the given latitude, which otherwise requires an extensive amount of snapshots to evaluate using simulations.

### v3

@@ III-A Invariant Probability Measure
Let $\overline{\mathcal{Q}}$ denote the product of two uniform distributions on the product space $S=[0,\frac{2\pi}{N_{o}}]\times[0,\frac{2\pi}{N_{s}}]$. Let $\overline{\mathcal{Q}}_{\Psi}$ denote the distribution of $\Psi$ at time zero when $(\bar{\theta},\bar{\omega})$ has distribution $\overline{\mathcal{Q}}$.
@@
The probability measure $\overline{\mathcal{Q}}$, on $S$, is invariant for $\{R_{t}\}$ and the probability measure $\overline{\mathcal{Q}}_{\Psi}$, on $\mathcal{N}_{r}$, is invariant for $\{\Psi_{t}\}$.
@@
It is enough to prove the first statement. Let $\mathbb{T}[0,a]\times[0,b]$ denote the torus on $[0,a]\times[0,b]$, with $a,b$ positive real numbers. Let $\theta=\bar{\theta}\frac{N_{o}}{2\pi}$ and $\omega=\bar{\omega}\frac{N_{s}}{2\pi}$. Let also $\theta_{t}=\bar{\theta}_{t}\frac{N_{o}}{2\pi}$ and $\omega_{t}=\bar{\omega}_{t}\frac{N_{s}}{2\pi}$. Denote by $g$ the linear map that transforms $(\theta_{t},\omega_{t})$ into $(\bar{\theta}_{t},\bar{\omega}_{t})$. The associated measurable flow is now $\alpha_{t}({\theta},{\omega})=({\theta}_{0}-v_{{\theta}}t\mod 1,{\omega}_{0}+v_{{\omega}}t\mod 1)$, where $v_{{\theta}}=\bar{v}_{{\theta}}\frac{N_{o}}{2\pi}$ and $v_{{\omega}}=\bar{v}_{{\omega}}\frac{N_{s}}{2\pi}$.
Then, $\overline{\mathcal{Q}}$ on $\bar{\mathbb{T}}([0,{2\pi}/N_{o}],[0,{2\pi}/N_{s}])$ is invariant for $\bar{\alpha}_{t}(\bar{\theta},\bar{\omega})=(\bar{\theta}-\bar{v}_{\theta}t{\mod 2\pi/N_{s}},\bar{\omega}+\bar{v}_{\omega}t{\mod 2\pi/N_{o}})$ iff $\mathcal{Q}$ on $\mathbb{T}([0,1],[0,1])$, namely the product of the two uniform distributions on [0,1][0,1][ 0 , 1 ], is left invariant by $\alpha_{t}(\theta,\omega)$. Then, $\overline{\mathcal{Q}}$ is left invariant by $\alpha_{t}(\theta,\omega)$ for all nonzero $(v_{{\theta}},v_{{\omega}})$ [12]. ∎
@@ III-B Ergodicity and Periodicity
If the ratio $\rho$ is irrational, then the dynamical system $\{R_{t}\}$ (resp. $\{\Psi_{t}\}$) on $S$ (resp. ${\mathcal{N}}_{r}$) is minimal and ergodic w.r.t. $\overline{\mathcal{Q}}$ (resp. ergodic w.r.t. ${\overline{\mathcal{Q}}}$). If the ratio $\rho$ is rational, then $\{R_{t}\}$ and $\{\Psi_{t}\}$ are both periodic.
@@
It is enough to prove the statement concerning $R$. We use the notation in the proof of Theorem 1. Let $\hat{\rho}$ denote the normalized ratio ${v_{\theta}}/{v_{\omega}}$. If $\hat{\rho}$ is irrational, then the dynamical system $\alpha_{t}$ on $\mathbb{T}([0,1],[0,1])$ is minimal and ergodic w.r.t. $\mathcal{Q}$, namely the product of two uniform distributions on [0,1][0,1][ 0 , 1 ] (see e.g. [13, Section 6.5.2] and [14, Theorem 2.1]). If $\hat{\rho}$ is rational (which is equivalent to having the ratio $\rho$ rational), then $\alpha_{t}$ is periodic on $\mathbb{T}([0,1],[0,1])$. Hence, it is not ergodic w.r.t. $\mathcal{Q}$ (see [13, Section 6.5.2] and [14, Theorem 2.1]). ∎
@@
In the irrational case, the constellation seen for a fixed Earth user is not periodic. Instead, we have
$\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}h(\Psi_{t}(x))\mathrm{d}t=\int h(\psi){\overline{\mathcal{Q}}}_{\Psi}(\mathrm{d}\psi),$ (8)
[…]
@@
Since the altitude basically determines the satellite speed, the designer can choose an altitude making the constellation periodic to fixed Earth observers. However, an unaware designer who would pick a real number, say between 350350350 km to 150015001500 km, would get an irrational ratio for almost all choices w.r.t. the Lebesgue measure on this interval since the rational numbers have Lebesgue measure zero in the given interval. Therefore, based on Theorems 1 and 2, one can claim that designers unaware of this dynamical system property have “almost surely” built constellations which are aperiodic and ergodic, and where the long time empirical averages converge to the expectation over the invariant distribution $\overline{\mathcal{Q}}_{\Psi}$ studied in the present paper.
@@
Assume that the designer selects the angular parameters $\bar{\theta}$ and $\bar{\omega}$ as random variables sampled independently, with each of them being uniform on the relevant interval. Then, when deconditioning the position of the satellites of the constellation w.r.t. these two random variables, one gets a distribution for the satellite positions which is invariant over time, and which is precisely the distribution ${\overline{\mathcal{Q}}}_{\Psi}$ of this paper and leads to the specific distributions that we evaluate below for the metrics of interest. By the strong law of large numbers, the distributions computed here can hence also be seen as the empirical averages of the distributions evaluated over a large collection of samples of the constellation obtained from i.i.d. samples of these two angles. This observation is valid in all cases including the periodic case. Of course, in the case of ergodicity, the metrics evaluated below are also time averages.
@@ IV-A Distance to Nearest Satellite
Let $\operatorname{\mathbb{S}}$ be the sphere of radius $r.$ For $r-e<d<\sqrt{r^{2}-e^{2}},$ we define the spherical cap w.r.t. the typical user $u$ and the distance $d$ as follows: $\operatorname{\mathbb{S}}_{u}(d)=\{(x,y,z)\in\operatorname{\mathbb{S}}|\|(x,y,z)-\vec{u}\|\leq d\}$. In particular, $\operatorname{\mathbb{S}}_{u}(\sqrt{r^{2}-e^{2}})$ is the visible spherical cap from the typical user and we simply denoted it by $\operatorname{\mathbb{S}}_{u}.$ Let $D(l_{u})$ be the distance from the typical user at latitude $l_{u}$ to its nearest satellite. We set $D(l_{u})=\infty$ if no satellite is visible from the typical user.
@@
$\zeta\int_{0}^{\frac{2\pi}{N_{o}}}\int_{0}^{\frac{2\pi}{N_{s}}}\!\!\mathbbm{1}\left\{\frac{r^{2}+e^{2}-d^{2}}{2re}>\max\frac{\vec{X}_{i,j}\cdot\vec{u}}{\|\vec{X}_{i,j}\|\|\vec{u}\|}\right\}\mathop{}\!\mathrm{d}\omega\mathop{}\!\mathrm{d}\theta,$
where $\vec{X}_{i,j}$ are in Eqs. (4)–(6), and $\zeta=N_{o}N_{s}/(4\pi^{2}).$
@@
Let $\mathbbm{1}\{x\}$ be the indicator function taking value one if $x$ occurs and zero otherwise. Conditionally on $\bar{\theta},\bar{\omega},$ the distance from the typical user to its nearest satellite is
@@
where $(i,j)\in([N_{o}],[N_{s}])$ stands for $i=1,...,N_{o}$ and $j=1,...,N_{s}.$ To get this, we use the fact that all points of such orbits are not in the cap if and only if the central angle between the user and the satellite closest to the user is greater than $\kappa,$ where $\kappa=\arccos\left(\frac{e^{2}+r^{2}-d^{2}}{2re}\right)$ is the central angle from the user to the rim of the cap $\operatorname{\mathbb{S}}_{u}(d)$. Then,
@@
$=\zeta\int_{0}^{\frac{2\pi}{N_{o}}}\int_{0}^{\frac{2\pi}{N_{s}}}\mathbbm{1}\left\{\cos(\kappa)>\max_{(i,j)}\frac{\vec{X}_{i,j}\cdot\vec{u}}{\|X_{i,j}\|\|u\|}\right\}\mathop{}\!\mathrm{d}\bar{\omega}\mathop{}\!\mathrm{d}\bar{\theta}$
$=\zeta\int_{0}^{\frac{2\pi}{N_{o}}}\!\int_{0}^{\frac{2\pi}{N_{s}}}\!\!\mathbbm{1}\left\{\frac{r^{2}+e^{2}-d^{2}}{2re}>\max_{(i,j)}\frac{\vec{X}_{i,j}\cdot\vec{u}}{\|\vec{X}_{i,j}\|\|\vec{u}\|}\right\}\!\mathop{}\!\mathrm{d}\bar{\omega}\mathop{}\!\mathrm{d}\bar{\theta},$
@@
Thanks to the deterministic structure of the deployment of satellites, for latitudes $l_{u}$ smaller than a threshold $L$, for $N_{o}$ and $N_{s}$ large enough, there exists a constant $d_{c}(l)\leq\sqrt{r^{2}-e^{2}}$ such that $\operatorname{\mathbf{P}}[r-e\leq D(l_{u})<d_{c}(l_{u})]=1$. For instance, when $N_{o}=30$ and $N_{s}=50,$ we have $\operatorname{\mathbf{P}}(r-e\leq D({l_{u}})<d_{c}(l_{u}))=1$ with $d_{c}(l)\approxeq 775$km, e.g., the association distance is for sure less than 775775775 km. This property is useful to establish performance lower bounds since it directly links to the maximum path loss of the downlink communications. The existence of this critical distance comes from the fact that proposed network features regularly placed satellites with a rigid structure, in contrast to other models based on binomial, Poisson, or Cox point processes [2, 3, 4, 5, 6, 7, 8]. Finding an explicit expression for $d_{c}(l)$ is left for future work.
@@
where the left-hand side is the time average of the association distance for a user and the right-hand side is its expectation w.r.t. the invariant distribution, that was obtained in Theorem 3.
@@
Fig. 2 shows the mean of the total interference. The simulation results confirm the accuracy of the derived formula. Note, the average total interference is the maximal for users at latitude 30°30$\degree$30 °.
@@
$\tilde{J}=\mkern 1.5mu\overline{\mkern-1.5muF\mkern-1.5mu}\mkern 1.5mu_{H}\left(\frac{\tau\|x_{\star}-u\|^{\alpha}}{pG_{\star}}\left(\sigma^{2}+\sum\!\!\frac{ph\tilde{c}_{i,j}G_{i,j}}{\|x_{i,j}-u\|^{\alpha}}\right)\right).$
@@
with $\tilde{c}_{i,j}\!=\!\mathbbm{1}\!\left\{\frac{\vec{X}_{i,j}\cdot\vec{u}}{\|\vec{X}_{i,j}\|\|\vec{u}\|}\geqq\frac{e}{r},\frac{\vec{X}_{i,j}\cdot\vec{u}}{\|\vec{X}_{i,j}\|\|\vec{u}\|}\neq\max\left(\frac{\vec{X}_{i,j}\cdot\vec{u}}{\|\vec{X}_{i,j}\|\|\vec{u}\|}\right)\!\right\}$,
@@
As in Theorem 4, Theorem 5 yields a simpler expression for a Rayleigh fading scenario where $\mkern 1.5mu\overline{\mkern-1.5muF\mkern-1.5mu}\mkern 1.5mu_{H}(x)=e^{-x}$. Fig. 3 illustrates the coverage probability for the Rayleigh fading. The figure confirms the accuracy of the derived formula.
The practical implication of employing the proposed geometric model lies in the fact that the derived formula computes the coverage probability of the typical user at a given latitude and it is equal to the coverage probability of all users at the same latitude. In addition, in the ergodic case of Theorem 2, the expression theoretically matches the empirically-obtained time average of the coverage probability of any user at the given latitude, which otherwise requires an extensive amount of snapshots to evaluate using simulations.